OpenStudy (anonymous):

intergrate cos ^2 (3X)

OpenStudy (psymon):

You have to use a power-reduction formula: \[\cos^{2}x = \frac{ 1+\cos2x }{ 2 }\]So basically, you rewrite your function as: \[\cos^{2}(3x) = \frac{1+\cos(6x)}{2}\]Then you can integrate like normal :3

OpenStudy (anonymous):

if we change it to \[1-\sin ^2 (3X)\]

OpenStudy (psymon):

That doesnt really make things any different. It would still be the same problem, except if you were to do a power-reduction formula with sin, itd be \[\sin^{2}x = \frac{1-\cos2x}{2}\] Theres no real use of using an identity other than the power-reduction one.

OpenStudy (anonymous):

ok thank you :)

OpenStudy (psymon):

mhm : )